OpenStudy (anonymous):

5x+y=0 5x+2y=30 Solve by elimination need help to make sure im right :D

OpenStudy (anonymous):

5x+y=0 5x=-y x=-y/5 plug this into the other equation... 5x+2y=30 5(-y/5)+2y=30 -y+2y=30 y=30 x=-30/5=-6

OpenStudy (anonymous):

but i need a little more explanation...like a walk through of all the steps

OpenStudy (anonymous):

Start with one equation: 5x+y=0 Subtract y from both sides: 5x=-y Divide both sides by 5: x=-y/5 Plug this value of x into the x in the other equation (5x+2y=30): 5(-y/5)+2y=30 Simplify:-y+2y=30 -y+2y=y, so: y=30 Then plug this back into x=-y/5 (from the other equation): x=-30/5 Simplify: x=-6 And that is the answer as well as I can explain it! :)

OpenStudy (anonymous):

thank you